NvidiaIntely (nvidiaintely):

A balloon containing 0.10 moles of gas has a volume of 2.0 liters. If temperature and pressure remain constant, what will be the new volume when 0.40 moles of gas are added to the balloon? 0.10 L 0.50 L 10.0 L 8.0 L

NvidiaIntely (nvidiaintely):

@Kevin @Will.H

OpenStudy (will.h):

heey

OpenStudy (will.h):

0.5 moles / 0.1 moles (2.0 L) = 10 L

OpenStudy (kevin):

From pv = nrt nrt/pv = nrt/pv n1/v1 = n2/v2 0.10/2 = (0.10+0.40)/v2 v2 = 10l

NvidiaIntely (nvidiaintely):

Thanks guys!